I need to convert this to Vertex form i dont know how y=-3x^2 +6x-5 please help me
Answer by scott8148(6628)

separate the constant term ___ y = (-3x^2 + 6x) - 5

factor out the squared term coefficient ___ y = -3(x^2 - 2x) - 5

complete the square ___ y = -3(x^2 - 2x + 1) - 5 + 3
___ remember to add a complementary term to the constant to maintain the equality

write in vertex form ___ y = -3(x - 1)^2 - 2